# Modal logics of almost sure validities in some classes of euclidean and transitive frames

### v2

@@ 1.2 Operations on frames
Let $X$ be a set. The diagonal relation on $X$ is $Id_{X}=\{(a,\,a)\mid a\in X\}.$ If $R\subseteq X\times X$ is a relation, let $R^{0}=Id_{X}$ and $R^{i+1}=R\circ R^{i}$ for any $i\in\omega.$
For any $U\subseteq X,$ $R[U]$ denotes the set $\{a\in X\mid\exists u\in U\>\left(uRa\right)\}.$ The notation $R(a)$, where $a\in X,$ abbreviates $R[\{a\}].$
@@
If $F=(X,\,R)$ is a frame and $U\subseteq X$, the subframe of $X$ generated by $U$ is the frame $F{\uparrow}U=(R^{*}[U],\,R{\upharpoonright}R^{*}[U]).$ If $a\in X$, then $F{\uparrow}a$ is a shorthand for $F{\uparrow}\{a\}.$ A frame $F$ is said to be point-generated if $F=F{\uparrow}a$ for some $a\in\operatorname{dom}F.$ The generated subframe preserves the validity of modal formulas: for any $U\subseteq\operatorname{dom}F$, $\operatorname{Log}{F}\subseteq\operatorname{Log}{F{\uparrow}U}$ [BdRV01, Proposition 2.6].
@@
Given a pair of frames $F=(X,\,R)$ and $G=(Y,\,S),$ a p-morphism from $F$ to $G$ is a surjective map $f:\>X\to Y$ such that $S(f(a))=f(R(a))$ for any $a\in X.$ We write $F\twoheadrightarrow G$ if there exists a p-morphism from $F$ to $G$. The p-morphism preserves the validity of modal formulas: if $F\twoheadrightarrow G$, then $\operatorname{Log}F\subseteq\operatorname{Log}G$ [BdRV01, Proposition 2.14]
@@ 1.3 Classes of frames and their logics
    serial if $\forall a\in X\>\left(R(a)\neq\varnothing\right);$
@@ 1.4 Random frames
By Theorem 1.1 $L^{\mathbf{as}}\subseteq(L^{\mathbf{as}})^{\mathbf{as}}.$ For the other direction, observe that $\operatorname{Fr}L^{\mathbf{as}}\subseteq\operatorname{Fr}L$ since $L\subseteq L^{\mathbf{as}}$, so for any $\varphi\in(L^{\mathbf{as}})^{\mathbf{as}}$,
@@
$=\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\cap\operatorname{Fr}\{\varphi\}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\right|}.$
$=\operatorname{P}(\hat{F}_{n}(L^{\mathbf{as}})\models\varphi)\cdot\operatorname{P}(\hat{F}_{n}(L)\in\operatorname{Fr}L^{\mathbf{as}})=1\cdot 1=1.$
@@ 2 Connected frames
Let $L$ be a modal logic such that $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}L$ is nonempty for any $n\in\omega.$ Then $L^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}L).$
@@
$\operatorname{P}(\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}L)\models\varphi)=\frac{|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}L\cap\operatorname{Fr}\{\varphi\}|}{|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}L|}\not\to 1,\,n\to\infty.$
@@ 3 Euclidean frames
Let $U=\bigcup_{a\in X}R(a).$ For any $u\in U,\,aRu$ for some $a\in X,$ so by the definition of Euclidean relation $uRu.$ Then $R{\upharpoonright}U$ is reflexive and Euclidean, hence an equivalence relation.
Every state $a\in X$ with $R{}^{-1}(a)\neq\varnothing$ is in $U$, so $R\subseteq X\times U.$
To prove the uniqueness, let us assume that $V\subseteq X$ satisfies the conditions. If $a\in V\setminus U$, then $R{}^{-1}(a)=\varnothing,$ so $R$ is not reflexive on $V\setminus U.$ Since $R{\upharpoonright}V$ is an equivalence relation, $V\subseteq U.$ If $a\in U$, then $aRa$, so $(a,\,a)\in X\times V,$ hence $a\in V.$ Then $U=V$, so $U$ is the unique subset of $X$ that satisfies the conditions.
@@
Conversely, suppose that $R{\upharpoonright}U_{F}$ is a cluster. For any $a\in X,$ $R(a)\subseteq U_{F}$ and $R(a)\neq\varnothing,$ so $aRu$ for some $u\in U.$ Then $aR^{2}v$ for any $v\in U.$ Therefore $R^{2}(X)=U,$ so $R^{2}\cup R^{-2}=X\times X,$ and $F$ is connected.
@@
A frame $F=([n],\,R)\in\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}$ is uniquely determined by $U_{F}$ and the family of subsets $\{R(a)\mid a\in[n]\setminus U_{F}\}$ where $R(a)\subseteq U_{F}$ and $R(a)\neq\varnothing$ for all $a.$ Then
@@
Let $\mathcal{Q}_{n,r}$ be the subset of $F=([n],\,R)\in\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}$ consisting of all frames $F=([n],R)$ such that for any $a\in[n]\setminus U_{F},$ ${|R(a)|\leq r}$ or $|R(a)|\geq|U_{F}|-r$. Let us estimate $|\mathcal{Q}_{n,r}|.$ If we fix $U_{F}\subseteq[n]$ with $|U_{F}|=m$, then for any $a\not\in U_{F}$ the nonempty subset $R(a)\subseteq U_{F}$ can be chosen in $\sum_{k=1}^{r}\left(\binom{m}{k}+\binom{m}{m-k}-1\right)$ ways. Note that
@@
$\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})=\mathbf{KD5}.$
@@
By Proposition 3.5 every connected Euclidean frame is serial, so $\hat{F}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})\models\mathbf{KD5}$ almost surely.
For the other direction, consider a formula $\varphi\not\in\mathbf{KD5}.$ There exists a finite point-generated $\mathbf{KD5}$-frame $G=(Y,S)=G{\uparrow}c$ such that $G\not\models\varphi.$ Let $r=|\operatorname{dom}G|.$ Let $F=(X,\,R)$ be a connected $\mathbf{KD5}$-frame such that
$|U_{F}|>r\text{ and }\exists a\in X\setminus U_{F}\>\left(r<|R(a)|<|U_{F}|-r\right).$ (13)
We define a p-morphism ${f:F{\uparrow}a\twoheadrightarrow G}.$ By Proposition 3.5, $\operatorname{dom}F{\uparrow}a=R^{*}(a)=U_{F}\cup\{a\}.$ Let $f(a)=c.$ Observe that $|R(a)|>r\geq|U_{G}|>|S(c)|,$ so let $f$ map $R(a)$ surjectively onto $S(c).$ Analogously, let $f{\upharpoonright}_{U_{F}\setminus R(a)}$ be a surjection onto $U_{G}\setminus S(c).$
@@
By Theorem 1.1, $\mathbf{KD5}\subseteq\mathbf{KD5}^{\mathbf{as}}$. For the other direction, observe that the cluster $([n],\,[n]\times[n])$ is a connected $\mathbf{KD5}$-frame for all $n\in\omega$, thus $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}\neq\varnothing$ for all $n\in\omega.$ By Theorem 2.9 $\mathbf{KD5}^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})=\mathbf{KD5}.$
@@
Let us find $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45}).$ Notice that a Euclidean frame $F=(X,R)$ is serial, transitive and connected iff $U_{F}\neq\varnothing$ and $R=X\times U_{F}$. Then such frame is uniquely determined by its cluster $U_{F}\neq\varnothing$. Therefore
@@
Then for any finite point-generated frame $G=(Y,\,S)\models\mathbf{KD45}$, a.a.s. there exist a p-morphism from some generated subframe of $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})$ to $G,$ so  $\operatorname{Log}\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\subseteq\operatorname{Log}G$ a.a.s., thus $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\subseteq\mathbf{KD45}.$
@@
The connected components of a $\mathbf{K5B}$-frame are clusters and irreflexive singletons. Then $\hat{F}_{n}(\mathbf{K5B})$ a.a.s. contains a cluster of size greater than $r$ for any fixed $r\in\omega$ by Proposition 2.6, so there is a p-morphism from a generated subframe of $\hat{F}_{n}$ to any finite cluster, so $\operatorname{Log}\hat{F}_{n}$ is contained in the logic of a finite cluster a.a.s.
A $\mathbf{K5B}$-frame with set of states $n$ is uniquely determined by a subset $E\subseteq[n]$ that consists of the irreflexive singletons and an equivalence relation on $n\setminus E.$ Then
@@
as $n\to\infty,$ so $\operatorname{Log}(\mathbf{K5B})$ is contained in the logic of an irreflexive singleton a.a.s.
Notice that $\mathbf{K5B}$ is the logic of its finite point-generated frames, which are exactly the finite clusters and the irreflexive singletons. Then $\mathbf{K5B}^{\mathbf{as}}\subseteq\mathbf{K5B}.$ The converse inclusion is true by Theorem 1.1.
@@
Observe that $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{S5}$ consists of one frame $(n,\,n\times n).$ The finite point-generated frames of $\mathbf{S5}$ are the finite clusters, which are p-morphic images of $(n,\,n\times n)$ for $n$ sufficiently large. Then $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{S5})=\mathbf{S5}.$ Then the statement of this theorem follows from Theorem 1.1 and Theorem 2.9.

### v3

@@ 1.2 Operations on frames
Let $X$ be a set. The diagonal relation on $X$ is $Id_{X}=\{(a,\,a)\mid a\in X\}.$ If $R\subseteq X\times X$ is a relation, let $R^{0}=Id_{X}$ and $R^{i+1}=R\circ R^{i}$ for any $i\in\omega.$ The inverse relation $R{}^{-1}$ is defined as $\{(a,b)\in X\mid bRa\},$ and $R^{-i}\equiv{(R{}^{-1})}^{i}$ for any $n\in\omega.$
For any $U\subseteq X,$ $R_{\mathrm{out}}[U]$ denotes the set $\{a\in X\mid\exists u\in U\>\left(uRa\right)\},$ and $R_{\mathrm{in}}[U]$ denotes $\{a\in X\mid\exists u\in U\>\left(aRu\right)\}.$ The notations $R_{\mathrm{out}}(a),\,R_{\mathrm{in}}(a)$, where $a\in X,$ abbreviate $R_{\mathrm{out}}[\{a\}]$ and $R_{\mathrm{in}}[\{a\}],$ respectively.
@@
If $F=(X,\,R)$ is a frame and $U\subseteq X$, the subframe of $X$ generated by $U$ is the frame $F{\uparrow}U=(R^{*}_{\mathrm{out}}[U],\,R{\upharpoonright}R^{*}_{\mathrm{out}}[U]).$ If $a\in X$, then $F{\uparrow}a$ is a shorthand for $F{\uparrow}\{a\}.$ A frame $F$ is said to be point-generated if $F=F{\uparrow}a$ for some $a\in\operatorname{dom}F.$ The generated subframe preserves the validity of modal formulas: for any $U\subseteq\operatorname{dom}F$, $\operatorname{Log}{F}\subseteq\operatorname{Log}{F{\uparrow}U}$ [BdRV01, Proposition 2.6].
@@
Given a pair of frames $F=(X,\,R)$ and $G=(Y,\,S),$ a p-morphism from $F$ to $G$ is a surjective map $f:\>X\to Y$ such that $S_{\mathrm{out}}(f(a))=f(R_{\mathrm{out}}(a))$ for any $a\in X.$ We write $F\twoheadrightarrow G$ if there exists a p-morphism from $F$ to $G$. The p-morphism preserves the validity of modal formulas: if $F\twoheadrightarrow G$, then $\operatorname{Log}F\subseteq\operatorname{Log}G$ [BdRV01, Proposition 2.14]
@@ 1.3 Classes of frames and their logics
    serial if $\forall a\in X\>\left(R_{\mathrm{out}}(a)\neq\varnothing\right);$
@@ 1.4 Random frames
By Theorem 1.1 $L^{\mathbf{as}}\subseteq(L^{\mathbf{as}})^{\mathbf{as}}.$ For the other direction, consider any $\varphi\in(L^{\mathbf{as}})^{\mathbf{as}}$. By the definition, there exist the limits:
$\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\cap\operatorname{Fr}\{\varphi\}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}=\operatorname{P}(\hat{F}_{n}(L^{\mathbf{as}})\models\varphi)=1,$
$\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\right|}=\operatorname{P}(\hat{F}_{n}(L)\models L^{\mathbf{as}})=1.$
Since $L\subseteq L^{\mathbf{as}}$, we have $\operatorname{Fr}L^{\mathbf{as}}\subseteq\operatorname{Fr}L$ , so for any $\varphi\in(L^{\mathbf{as}})^{\mathbf{as}}$,
@@
$=\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\cap\operatorname{Fr}\{\varphi\}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}\lim_{n\to\infty}\frac{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L^{\mathbf{as}}\right|}{\left|\mathcal{F}_{n}\cap\operatorname{Fr}L\right|}=1\cdot 1=1.$
@@ 2 Connected frames
Let $L$ be a modal logic such that:
[…]
    $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}L$ is nonempty for any $n\in\omega.$
[…]
    For any $\varphi\not\in\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}L),$
    $\limsup_{n\to\infty}\operatorname{P}(\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}L)\models\varphi)<1.$ (9)
Then $L^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}L).$
@@
$\limsup_{n\to\infty}\operatorname{P}(\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}L)\models\varphi)=\limsup_{n\to\infty}\frac{|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}L\cap\operatorname{Fr}\{\varphi\}|}{|\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}L|}<1.$
@@ 3 Euclidean frames
Let $U=\bigcup_{a\in X}R_{\mathrm{out}}(a).$ For any $u\in U,\,aRu$ for some $a\in X,$ so by the definition of Euclidean relation $uRu.$ Then $R{\upharpoonright}U$ is reflexive and Euclidean, hence an equivalence relation.
Every state $a\in X$ with $R_{\mathrm{in}}(a)\neq\varnothing$ is in $U$, so $R\subseteq X\times U.$
To prove the uniqueness, let us assume that $V\subseteq X$ satisfies the conditions. If $a\in V\setminus U$, then $R_{\mathrm{in}}(a)=\varnothing,$ so $R$ is not reflexive on $V\setminus U.$ Since $R{\upharpoonright}V$ is an equivalence relation, $V\subseteq U.$ If $a\in U$, then $aRa$, so $(a,\,a)\in X\times V,$ hence $a\in V.$ Then $U=V$, so $U$ is the unique subset of $X$ that satisfies the conditions.
@@
Conversely, suppose that $R{\upharpoonright}U_{F}$ is a cluster. For any $a\in X,$ $R_{\mathrm{out}}(a)\subseteq U_{F}$ and $R_{\mathrm{out}}(a)\neq\varnothing,$ so $aRu$ for some $u\in U.$ Then $aR^{2}v$ for any $v\in U.$ Therefore $R^{2}_{\mathrm{out}}[X]=U,$ so $R^{2}\cup R^{-2}=X\times X,$ and $F$ is connected.
@@
A frame $F=([n],\,R)\in\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}$ is uniquely determined by $U_{F}$ and the family of subsets $\{R_{\mathrm{out}}(a)\mid a\in[n]\setminus U_{F}\}$ where $R_{\mathrm{out}}(a)\subseteq U_{F}$ and $R_{\mathrm{out}}(a)\neq\varnothing$ for all $a.$ Then
@@
Let $\mathcal{Q}_{n,r}$ be the subset of $F=([n],\,R)\in\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{KD5}$ consisting of all frames $F=([n],R)$ such that for any $a\in[n]\setminus U_{F},$ ${|R_{\mathrm{out}}(a)|\leq r}$ or $|R_{\mathrm{out}}(a)|\geq|U_{F}|-r$. Let us estimate $|\mathcal{Q}_{n,r}|.$ If we fix $U_{F}\subseteq[n]$ with $|U_{F}|=m$, then for any $a\not\in U_{F}$ the nonempty subset $R_{\mathrm{out}}(a)\subseteq U_{F}$ can be chosen in $\sum_{k=1}^{r}\left(\binom{m}{k}+\binom{m}{m-k}-1\right)$ ways. Note that
@@
For any $\varphi\not\in\mathbf{KD5},\,\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD5})\not\models\varphi$ a.a.s.
@@
Consider a formula $\varphi\not\in\mathbf{KD5}.$ There exists a finite point-generated $\mathbf{KD5}$-frame $G=(Y,S)=G{\uparrow}c$ such that $G\not\models\varphi.$ Let $r=|\operatorname{dom}G|.$ Let $F=(X,\,R)$ be a connected $\mathbf{KD5}$-frame such that
$|U_{F}|>r\text{ and }\exists a\in X\setminus U_{F}\>\left(r<|R_{\mathrm{out}}(a)|<|U_{F}|-r\right).$ (14)
We define a p-morphism ${f:F{\uparrow}a\twoheadrightarrow G}.$ By Proposition 3.5, $\operatorname{dom}F{\uparrow}a=R^{*}_{\mathrm{out}}(a)=U_{F}\cup\{a\}.$ Let $f(a)=c.$ Observe that $|R_{\mathrm{out}}(a)|>r\geq|U_{G}|>|S(c)|,$ so let $f$ map $R_{\mathrm{out}}(a)$ surjectively onto $S(c).$ Analogously, let $f{\upharpoonright}_{U_{F}\setminus R_{\mathrm{out}}(a)}$ be a surjection onto $U_{G}\setminus S(c).$
@@
[…]
@@
Let $\varphi\not\in\mathbf{KD45},$ then $G\not\models\varphi$ for some finite point-generated frame $G\models\mathbf{KD45}$. Let us show that $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\twoheadrightarrow G$ a.a.s. Notice that a Euclidean frame $F=(X,R)$ is serial, transitive and connected iff $U_{F}\neq\varnothing$ and $R=X\times U_{F}$. Then such frame is uniquely determined by its cluster $U_{F}\neq\varnothing$. Therefore
@@
There a.a.s. exists a p-morphism from some generated subframe of $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})$ to $G,$ so $\operatorname{P}(\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\models\varphi)\to 0$ as $n\to\infty.$
Since $\varphi\not\in\mathbf{KD45}$ was arbitrary, $\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{KD45})\subseteq\mathbf{KD45}.$
@@
Notice that $\mathbf{K5B}$ is the logic of its finite point-generated frames, which are exactly the finite clusters and the irreflexive singletons. If $\varphi\not\in\mathbf{K5B}$ for some formula $\varphi,$ then either $([r],\,[r]\times[r])\not\models\varphi$ for some $r\in\omega$, or ($(\{a\},\varnothing)\not\models\varphi.$
Let $r\in\omega$ and $([r],\,[r]\times[r])\not\models\varphi$. The connected components of a $\mathbf{K5B}$-frame are clusters and irreflexive singletons. Then $\hat{F}_{n}(\mathbf{K5B})$ a.a.s. contains a cluster of size greater than $r$ by Proposition 2.6, so there is a p-morphism from a generated subframe of $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5B}))$ to $([r],[r]\times[r])$, so $\operatorname{P}(\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5B})\models\varphi)\to 0$ as $n\to\infty.$
Next we consider a formula $\varphi$ such that ($(\{a\},\varnothing\not\models\varphi)$. A $\mathbf{K5B}$-frame with set of states $n$ is uniquely determined by a subset $E\subseteq[n]$ that consists of the irreflexive singletons and an equivalence relation on $n\setminus E.$ Then
@@
as $n\to\infty.$ Then $\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5B})$ contains a generated subframe isomorphic to ($(\{a\},\varnothing)$ a.a.s., so $\operatorname{P}(\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{K5B})\models\varphi)\to 0$ as $n\to\infty.$
By Theorem 2.6, $\mathbf{K5B}^{\mathbf{as}}\subseteq\operatorname{Log}^{\mathbf{as}}(\operatorname{Con}\operatorname{Fr}\mathbf{K5B})\subseteq\mathbf{K5B}.$ The converse inclusion is true by Theorem 1.1.
@@
Observe that $\mathcal{F}_{n}\cap\operatorname{Con}\operatorname{Fr}\mathbf{S5}$ consists of one frame $(n,\,n\times n).$ The finite point-generated frames of $\mathbf{S5}$ are the finite clusters, which are p-morphic images of $(n,\,n\times n)$ for $n$ sufficiently large. Then $\operatorname{P}(\hat{F}_{n}(\operatorname{Con}\operatorname{Fr}\mathbf{S5})\models\varphi)\to 0$ for any $\varphi\not\in\mathbf{S5}.$ Then the statement of this theorem follows from Theorem 1.1 and Theorem 2.9.